I = 1000 PM = 20 PP = 5 • Where: I: income; PM: price of meat; PP: price of potato

1) Budget equation? Graph?

2) TU = ( M – 2 ) PM: quantity of meat, P: quantity of potato Which consumption possibility to maximize TU?

3) Write new budget equation when PP = 10. Which consumption possibility to maximize TU?


1
Expert's answer
2021-11-25T10:24:00-0500

1)Budget equation:

1000=20x+5y




2)

x=M, y=P

Product derivative x:

(TU = ( x – 2 ) y)'=y

Product derivative y:

(TU = ( x – 2 ) y)'=(x-2)

The consumer must spend his income in such a way that the ratio of marginal utility to price is the same for all goods, while the income must be spent completely:

"\frac{MUx}{Px} = \frac{MUy}{Py}"

"\frac{y}{20} = \frac{(x-2)}{5}"

y+8=4x

y=8;4 x=4;3

3)Budget equation:

1000=20x+10y

"\frac{y}{20} = \frac{(x-2)}{10}"

y+4=2x

y=4; 2

x=3; ,2